5. An airplane accelerates down a runway at 3.7 m/s2 for 30.8s until it finally lifts off the ground. Determine the distance traveled before takeoff.
1
Expert's answer
2020-09-21T08:13:12-0400

A distance trevelled by a body with a constant acceleration (starting from the rest) is given by the following expression:


"d = \dfrac{at^2}{2}"

where "a = 3.7m\/s^2" is the acceleration and "t = 30.8s" is the time of trevelling. Thus, obtain:


"d = \dfrac{at^2}{2} = \dfrac{3.7\cdot 30.8^2}{2} =1754.984m"

Answer. 1754.984 m.